# Lepton Number Violation Higgs Decay at Muon Collider

### v1

@@ IV Signature from SM Higgs
where we have assumed that the heavy neutral lepton $N$ couples exclusively with the muon flavor leptons for illustration, i.e., $|V_{\mu N}|\neq 0$, and $|V_{eN}|=|V_{\tau N}|=0$. At the TeV scale muon collider, the dijets from heavy neutral lepton decays could be highly boosted and could merge into one fat-jet $J$. Hence, the signature is regarded as $\mu^{\pm}\mu^{\pm}JJ$. With narrow width approximation, the cross section of lepton number violation signature from SM Higgs decay can be calculated as
@@
The contributions of ${\mu^{+}}{\mu^{-}}\rightarrow{\mu^{+}}{\mu^{-}}WW$ and ${\mu^{+}}{\mu^{-}}\rightarrow{\mu^{+}}{\mu^{-}}WWZ$ are from lepton charge misidentified, which are significantly suppressed by the misidentification rate $0.1\%$ Liu:2021akf . Backgrounds from the ${\mu^{+}}{\mu^{-}}\rightarrow\mu^{\pm}{\nu}WWW$ process is dominant by the vector boson fusion, where the two same-sign $W$ bosons decay leptonically. Other backgrounds from $\mu^{+}\mu^{+}\to\mu^{+}\mu^{-}jj$, ${\mu^{+}}{\mu^{-}}\rightarrow{W^{\pm}}{W^{\pm}}jj$ and ${\mu^{+}}{\mu^{-}}\rightarrow t\bar{t}{W^{\pm}}{\mu^{\mp}}v$ are not considered, as they can be greatly reduced by proper cuts He:2024dwh .
In the simulation, we use FeynRules package Arun:2022ecj to implement the scalar singlet extension of the type-I seesaw model. We then use Madgraph5_aMC@NLO Alwall:2014hca to generate the signal and background events at the parton level. After this, Pythia8 Sjostrand:2014zea is used to do parton showering and hadronization, and the detector simulation is performed by Delphes3 deFavereau:2013fsa with the corresponding muon collider card. The fat-jets are reconstructed by using the Valencia algorithm Boronat:2014hva with $R=1.2$.
@@
$|\eta(\mu)|<2.5,P_{T}(J)>50~{}{\rm{GeV}},|\eta(J)|<2.5.$ (33)
In Figure 5, we show the distributions of some variables for the signal and backgrounds at the 3 TeV muon collider after the pre-selection cuts, where $m_{N}=25$GeV and 45GeV with $\sin\alpha=0.2$ are chosen as the benchmark points. Distributions of these variables at the 10 TeV muon collider are similar to the 3 TeV muon collider, hence they will not be shown in this paper.
@@
To make sure the same-sign dimuon come from the heavy neutral lepton, cuts on fat-jets are required. It is clear in Figure 5 that there are no peak structures around $m_{W}$ in the distribution of invariant mass of fat-jet $m_{J}$ for the signal, which conforms to the fact that these fat-jets are from the off-shell $W$ boson decay. As the same-sign dimuon cuts are already enough to suppress the SM backgrounds much smaller than the signal, we only require at least one fat-jet $J$ in the final states to keep as many signal events as possible. Based on the distribution of $P_{T}(J)$ in Figure 5, we also require the transverse momentum of fat-jets to be less than 200 GeV. The fat-jet cuts are
@@
where $N_{S}$ and $N_{B}$ are the event number of signal and backgrounds, respectively. To reach the $5\sigma$ discovery limit for the benchmark point, the required luminosity is about 21(4.4) fb-1 at the 3(10) TeV muon collider. Therefore, the lepton number violation SM Higgs decay signature is quite promising at the TeV scale muon collider.
Figure 6: Left panel: Significance for the lepton number violation SM Higgs decay $h\to\mu^{\pm}\mu^{\pm}JJ$. The blue, green, and red lines are the results with $\sin{\alpha}=0.2,0.15$, and 0.1 respectively. The dashed (solid) lines are for 3 (10) TeV muon collider with an integrated luminosity of $L=1(10)~{}\text{ab}^{-1}$. Right panel: The $5\sigma$ sensitivity reach of the $h\to\mu^{\pm}\mu^{\pm}JJ$ signature at the 3 TeV (dashed line) and 10 TeV (solid line) muon collider with an integrated luminosity of $L=1~{}\text{ab}^{-1}$ and $L=10~{}\text{ab}^{-1}$, respectively.
The signature of lepton number violation SM Higgs decay $h\to\mu^{\pm}\mu^{\pm}JJ$ depends on the mass of heavy neutral lepton $m_{N}$ and the Higgs mixing parameter $\sin\alpha$. We then perform a cut-based analysis to explore the sensitive region. In the left panel of Figure 6, we show the significance of the $h\to{\mu^{\pm}}{\mu^{\pm}}JJ$ signature by varying the mass of heavy neutral leptons $m_{N}$ for three values of mixing angle $\sin\alpha=0.2,0.15$ and 0.1 at the muon collider. With an integrated luminosity of $L=1(10)~{}\text{ab}^{-1}$, the $h\to{\mu^{\pm}}{\mu^{\pm}}JJ$ signature can be discovered in the mass range of $10(5)~{}\text{GeV}<m_{N}<60~{}\text{GeV}$ for $\sin\alpha=0.2$ at the 3 (10) TeV muon collider. The largest significance could reach about 30(230) when $m_{N}\simeq 45$ GeV for the collision energy of 3 (10) TeV. Lowing the mixing angle will lead to a smaller discovery mass range. The 3 TeV muon collider is promising to cover the mass range of $16~{}\text{GeV}<m_{N}<56~{}\text{GeV}$ when $\sin\alpha>0.1$. In the right panel of Figure 6, we depict the $5\sigma$ discovery reach of the $h\to\mu^{\pm}\mu^{\pm}JJ$ signature at the 3(10) TeV muon collider with an integrated luminosity of $L=1(10)~{}\text{ab}^{-1}$. When $m_{N}\simeq 45~{}\rm{GeV}$, the limit of the mixing parameter $\sin\alpha$ can be as small as $\sin\alpha=0.05(0.009)$ at the 3(10) TeV muon collider, which corresponds to the branching ratio of $h\to NN$ as $4.1\times 10^{-3}(1.3\times 10^{-4})$.
@@ V Signature from Heavy Higgs
For difference, we consider that the heavy neutral lepton is heavier than the $W$ boson. So the SM Higgs decay $h\to NN$ is forbidden in this scenario. Different from the signature $h\to\mu^{\pm}\mu^{\pm}JJ$ with fat-jets from the off-shell $W$ decay, the fat-jets in the signature $H\to\mu^{\pm}\mu^{\pm}JJ$ are from the on-shell $W$ decay when $m_{N}>m_{W}$. Hence, we expected the invariant mass of fat-jet satisfying $m_{J}\sim m_{W}$ in this scenario, which can be used to distinguish the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature from the $h\to\mu^{\pm}\mu^{\pm}JJ$ signal.
@@
[…]
Another important parameter that alerts the signal cross section is the Higgs mixing angle $\alpha$. In panel (c) of Figure 7, we show the cross section of $H\to\mu^{\pm}\mu^{\pm}JJ$ by varying $\sin\alpha$. Although the production cross section of heavy Higgs $\sigma(\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}H)$ decreases as $\sin\alpha$ becomes smaller according to Equation 38, the branching ratio of $H\to NN$ increases when $\sin\alpha$ decreases as shown in Figure 2. Therefore, the signal cross section $\sigma(H\to\mu^{\pm}\mu^{\pm}JJ)$ is nearly a constant when $\sin\alpha>0.1$. On the other hand, for $\sin\alpha<0.05$, the branching ratio BR$(H\to NN)$ is already close to one, so the signal cross section becomes proportional to $\sin^{2}\alpha$, and decreases quickly as $\sin\alpha$ becomes smaller.
@@
Figure 9: Upper panels: significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature. In the up-left panel we have fixed the mass relation $m_{N}=m_{H}/3$. The blue, green, and red lines are the results for $\sin{\alpha}=0.15,0.1$, and $0.05$, respectively. In the up-right panel we have fixed $\sin{\alpha}=0.1$. The blue, green, and red lines are the results for $m_{H}=$300 GeV, 500 GeV, 700GeV. The dashed (solid) lines are for 3 (10) TeV muon collider with an integrated luminosity of $L=1(10)~{}\text{ab}^{-1}$. Lower panels: the discovery reach of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature. In the down-left panel we have set $m_{N}=m_{H}/3$. The red region is excluded by direct searches of heavy Higgs Lane:2024vur . In the down-right panel, the red and green lines are the results for $\sin\alpha=0.1$ and $0.05$.
@@
Based on the selection cuts in Table 2, we then explore the significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature at the muon collider. The results are shown in Figure 9. In the left-upper panel, we show the significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature as a function of $m_{N}$ for $\sin\alpha=0.15,0.10$, and 0.05 respectively, where we have fixed the mass relation $m_{N}=m_{H}/3$. The significance decreases as $m_{N}$ increases. On the 3TeV muon collider with $1~{}\text{ab}^{-1}$ integrated luminosity, we can reach 5 $\sigma$ discovery for $m_{N}\lesssim 330$ GeV and $\sin\alpha=0.15$. Lowering $\sin\alpha$ to 0.05, the significance still exceeds 5 $\sigma$ when $m_{N}\lesssim 200$ GeV. Increasing the collision energy to 10 TeV with 10 ab-1 data could extend the discovery limit up to $m_{N}\lesssim 1$ TeV with $\sin\alpha=0.1$.
@@
In the right-upper panel of Figure 9, we show the significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature as a function of $m_{N}$ for $m_{H}=$300 GeV, 500 GeV and 700 GeV respectively, where we have fixed the Higgs mixing $\sin\alpha=0.1$. For $m_{H}=300$ GeV, the sensitivity reaches the maximum value of 23 $\sigma$ around $m_{N}=110$ GeV at the 3 TeV muon collider, and such result can reach to 130 $\sigma$ around $m_{N}=110$ GeV at the 10 TeV muon collider,. Meanwhile, for $m_{H}=500$ GeV and 700 GeV, we find that there exists a local minimum value of significance around $m_{N}\sim 135$ GeV. And the 5 $\sigma$ discovery could be realized when $m_{N}\lesssim 230$ GeV with $m_{H}=500$ GeV or $m_{N}\lesssim 320$ GeV with $m_{H}=700$ GeV at the 3 TeV muon collider, and $m_{N}\lesssim 240$ GeV with $m_{H}=500$ GeV or $m_{N}\lesssim 340$ GeV with $m_{H}=700$ GeV at the 10 TeV muon collider. Therefore, the 10 TeV muon collider is expected to probe most regions below the kinematic threshold $m_{N}\lesssim m_{H}/2$.
@@
The $5\sigma$ discovery region in the $m_{N}-\sin\alpha$ plane with $m_{N}=m_{H}/3$ is shown in the down-left panel of Figure 9. At the 3 TeV muon collider with $1~{}\text{ab}^{-1}$ data, the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature can be detected with the Higgs mixing parameter as small as $\sin\alpha\simeq 0.03$. According to the results in Figure 7, the cross section of $H\to\mu^{\pm}\mu^{\pm}JJ$ is heavily suppressed when $\sin\alpha\lesssim 0.01$, thus it becomes less promising. Due to the larger cross section and integrated luminosity, the 10 TeV muon collider could discover the Higgs mixing parameter as small as $\sin\alpha\simeq 0.005$ for heavy Higgs around the electroweak scale. With larger singlet scalar mass, the value of Higgs mixing parameter $\sin\alpha$ increases for 5 $\sigma$ discovery. Typically, for $\sin\alpha=0.15$, we might probe $m_{N}\lesssim 330$ GeV and $m_{N}\lesssim 950$ GeV for 3 TeV muon collider and 10 TeV muon collider separately.
@@
In the down-right panel of Figure 9, we show the sensitivity region in the $m_{H}-m_{N}$ plane with fixed $\sin\alpha$ of 0.1 and 0.05. At the 3 TeV muon collider, a large parameter space can be detected for $230~{}{\rm{GeV}}<m_{H}<880~{}\rm{GeV}$ and $m_{N}<340~{}\rm{GeV}$ with $\sin\alpha=0.1$. Of course, the sensitive region heavily depends on the Higgs mixing parameter. For $\sin\alpha>0.1$, although certain region with $m_{H}\lesssim 600$ GeV is already excluded by current direct searches of heavy Higgs, $m_{H}$ around TeV scale is also promising with such large $\sin\alpha$. When $\sin\alpha=0.05$, the promising region becomes $230~{}{\rm{GeV}}<m_{H}<640~{}\rm{GeV}$ and $m_{N}<220~{}\rm{GeV}$, which is clearly smaller than the scenario with $\sin\alpha=0.1$. Increasing the collision energy to 10 TeV will greatly enlarge the discovery region. For instance, the parameter space can be detected for $220~{}{\rm{GeV}}<m_{H}<2800~{}\rm{GeV}$ and $m_{N}<1100~{}\rm{GeV}$ with $\sin\alpha=0.1$, as well as $220~{}{\rm{GeV}}<m_{H}<2000~{}\rm{GeV}$ and $m_{N}<800~{}\rm{GeV}$ for $\sin\alpha=0.05$.

### v2

@@ IV Signature from SM Higgs
where we have assumed that the heavy neutral lepton $N$ couples exclusively with the muon flavor leptons for illustration, i.e., $|V_{\mu N}|\neq 0$, and $|V_{eN}|=|V_{\tau N}|=0$. At the TeV scale muon collider, the dijets from heavy neutral lepton decays could be highly boosted and could merge into one fat-jet $J$. Hence, the signature is regarded as $\mu^{\pm}\mu^{\pm}JJ$. For the electron mixing pattern, i.e., $|V_{eN}|\neq 0$, and $|V_{\mu N}|=|V_{\tau N}|=0$, the cross section of the signal will be the same as the muon mixing pattern. However, the cross sections of the backgrounds involving electron are usually smaller than those involving muon Li:2023tbx , so the electron mixing pattern is expected to be more promising than the muon mixing pattern at the muon collider. The muon mixing pattern in this paper represents a more conservative benchmark.
With narrow width approximation, the cross section of lepton number violation signature from SM Higgs decay can be calculated as
@@
The contributions of ${\mu^{+}}{\mu^{-}}\rightarrow{\mu^{+}}{\mu^{-}}WW$ and ${\mu^{+}}{\mu^{-}}\rightarrow{\mu^{+}}{\mu^{-}}WWZ$ are from lepton charge misidentified, which are significantly suppressed by the misidentification rate $0.1\%$ Liu:2021akf ; ATLAS:2019jvq . Backgrounds from the ${\mu^{+}}{\mu^{-}}\rightarrow\mu^{\pm}{\nu}WWW$ process is dominant by the vector boson fusion, where the two same-sign $W$ bosons decay leptonically. Other backgrounds from $\mu^{+}\mu^{+}\to\mu^{+}\mu^{-}jj$, ${\mu^{+}}{\mu^{-}}\rightarrow{W^{\pm}}{W^{\pm}}jj$ and ${\mu^{+}}{\mu^{-}}\rightarrow t\bar{t}{W^{\pm}}{\mu^{\mp}}v$ are not considered, as they can be greatly reduced by proper cuts He:2024dwh . For example, after the selection cuts in Table 2, the total cross section of these additional backgrounds is $8.4\times 10^{-5}$ fb at the 3 TeV muon collider, which is over two orders of magnitude smaller than the dominant backgrounds. So including these additional backgrounds has a negligible impact on the final results.
In the simulation, we use FeynRules package Alloul:2013bka to implement the scalar singlet extension of the type-I seesaw model. We then use Madgraph5_aMC@NLO Alwall:2014hca to generate the leading order signal and background events at the parton level. After this, Pythia8 Sjostrand:2014zea is used to do parton showering and hadronization, and the detector simulation is performed by Delphes3 deFavereau:2013fsa with the corresponding muon collider card. The fat-jets are reconstructed by using the Valencia algorithm Boronat:2014hva with $R=1.2$.
@@
$|\eta(\mu)|<2.5,P_{T}(J)>50~{\rm{GeV}},|\eta(J)|<2.5,$ (33)
which correspond to the default set of Delphes3.
In Figure 5, we show the distributions of some variables for the signal and backgrounds at the 3 TeV muon collider after the pre-selection cuts, where $m_{N}=25$GeV and 45GeV with $\sin\alpha=0.2$ are chosen as the benchmark points. Distributions of these variables at the 10 TeV muon collider are similar to the 3 TeV muon collider, hence they will not be shown in this paper. The distributions are normalized by using number of event in each bin divided by the total events. The normalized distributions are used to compare the difference between the signal and the backgrounds at the same level, based on which further selection cuts are applied.
@@
To make sure the same-sign dimuon come from the heavy neutral lepton, cuts on fat-jets are required. It is clear in Figure 5 that there are no peak structures around $m_{W}$ in the distribution of invariant mass of fat-jet $m_{J}$ for the signal, which conforms to the fact that these fat-jets are from the off-shell $W$ boson decay. The distributions of $m_{J}$ for the signal depend on the value of $m_{N}$. For heavier $m_{N}$, e.g., $m_{N}=45$ GeV, the distribution of $m_{J}$ will have a larger overlap with the background. We do not apply any cuts on $m_{J}$ in this analysis for simplicity. So the results in Table 1 might be improved by adopting $m_{N}$ dependent cut on $m_{J}$, e.g., $m_{J}<30$ GeV when $m_{N}=25$ GeV. As the same-sign dimuon cuts are already enough to suppress the SM backgrounds much smaller than the signal, we only require at least one fat-jet $J$ in the final states to keep as many signal events as possible. Based on the distribution of $P_{T}(J)$ in Figure 5, we also require the transverse momentum of fat-jets to be less than 200 GeV. The fat-jet cuts are
@@
where $N_{S}$ and $N_{B}$ are the event number of signal and backgrounds, respectively. To reach the $5\sigma$ discovery limit for the benchmark point, the required luminosity is about 21(4.4) fb-1 at the 3(10) TeV muon collider. Therefore, the lepton number violation SM Higgs decay signature is quite promising at the TeV scale muon collider. We then estimate the precision for the signature by using $\sqrt{N_{S}+N_{B}}/N_{S}$ Forslund:2022xjq . For the benchmark point in Table 1, we obtain a precision of 8.8% at the 3 TeV and 1.7% at the 10 TeV muon collider after all selection cuts.
Figure 6: Significance (upper-left panel) and precision (upper-right panel) for the lepton number violation SM Higgs decay $h\to\mu^{\pm}\mu^{\pm}JJ$. The blue, green, and red lines are the results with $\sin{\alpha}=0.2,0.15$, and 0.1 respectively. The dashed (solid) lines are for 3 (10) TeV muon collider with an integrated luminosity of $L=1(10)~\text{ab}^{-1}$. Lower panel: The $5\sigma$ sensitivity reach of the $h\to\mu^{\pm}\mu^{\pm}JJ$ signature at the 3 TeV (dashed line) and 10 TeV (solid line) muon collider with an integrated luminosity of $L=1~\text{ab}^{-1}$ and $L=10~\text{ab}^{-1}$, respectively.
The signature of lepton number violation SM Higgs decay $h\to\mu^{\pm}\mu^{\pm}JJ$ depends on the mass of heavy neutral lepton $m_{N}$ and the Higgs mixing parameter $\sin\alpha$. We then perform a cut-based analysis to explore the sensitive region. In the upper-left panel of Figure 6, we show the significance of the $h\to{\mu^{\pm}}{\mu^{\pm}}JJ$ signature by varying the mass of heavy neutral leptons $m_{N}$ for three values of mixing angle $\sin\alpha=0.2,0.15$ and 0.1 at the muon collider. With an integrated luminosity of $L=1(10)~\text{ab}^{-1}$, the $h\to{\mu^{\pm}}{\mu^{\pm}}JJ$ signature can be discovered in the mass range of $10(5)~\text{GeV}<m_{N}<60~\text{GeV}$ for $\sin\alpha=0.2$ at the 3 (10) TeV muon collider. For $m_{N}$ below 10 GeV, the final state muon from boosted $N$ decay can hardly be separated from the jets, which leads to the quickly decrease of the significance. The largest significance could reach about 30(230) when $m_{N}\simeq 45$ GeV for the collision energy of 3 (10) TeV. A smaller mixing angle will lead to a smaller discovery mass range. The 3 TeV muon collider is promising to cover the mass range of $16~\text{GeV}<m_{N}<56~\text{GeV}$ when $\sin\alpha>0.1$. The precision results are shown in the upper-right panel of Figure 6. Typically, the 3 TeV muon collider could have a precision of $\mathcal{O}(10)$% for $\sin\alpha\gtrsim 0.1$. And the 10 TeV muon collider could reach the precision of $\mathcal{O}(2)$% for $\sin\alpha\simeq 0.2$ with 25 GeV $\lesssim m_{N}\lesssim$ 55 GeV.
In the lower panel of Figure 6, we depict the $5\sigma$ discovery reach of the $h\to\mu^{\pm}\mu^{\pm}JJ$ signature at the 3(10) TeV muon collider with an integrated luminosity of $L=1(10)~\text{ab}^{-1}$. When $m_{N}\simeq 45~\rm{GeV}$, the limit of the mixing parameter $\sin\alpha$ can be as small as $\sin\alpha=0.05(0.009)$ at the 3(10) TeV muon collider, which corresponds to the branching ratio of $h\to NN$ as $4.1\times 10^{-3}(1.3\times 10^{-4})$.
@@ V Signature from Heavy Higgs
For the sake of difference, we consider that the heavy neutral lepton is heavier than the $W$ boson. So the SM Higgs decay $h\to NN$ is forbidden in this scenario. Different from the signature $h\to\mu^{\pm}\mu^{\pm}JJ$ with fat-jets from the off-shell $W$ decay, the fat-jets in the signature $H\to\mu^{\pm}\mu^{\pm}JJ$ are from the on-shell $W$ decay when $m_{N}>m_{W}$. Hence, we expected the invariant mass of fat-jet satisfying $m_{J}\sim m_{W}$ in this scenario, which can be used to distinguish the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature from the $h\to\mu^{\pm}\mu^{\pm}JJ$ signal.
@@
$\sigma(\mu^{+}\mu^{-}\xrightarrow{H}\mu^{\pm}\mu^{\pm}JJ)\simeq\sigma(\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}H)\times\text{BR}(H\to NN)\times\text{BR}(N\to\mu^{\pm}jj)^{2}/2,$ (40)
In Figure 7, we show the numerical results of $\sigma(H\to\mu^{\pm}\mu^{\pm}JJ)$ at the muon collider. In panel (a), the cross section at 3 TeV and 10 TeV muon collider as a function of $m_{H}$ are shown, where we set $m_{N}=m_{H}/3$ for illustration. We should mention that for $\sin\alpha=0.15$, the current direct searches of heavy Higgs boson could exclude the mass region of 300 $\lesssim m_{H}\lesssim 500 GeV$, meanwhile, all the mass range is allowed for $\sin\alpha<0.1$ Lane:2024vur . The cross section decreases as the heavy Higgs mass increases. Typically, we have $\sigma(H\to\mu^{\pm}\mu^{\pm}JJ)>0.01$ fb when $m_{H}\lesssim 1.5(5)$ TeV at the 3 (10) TeV muon collider. Hence, increasing the collision energy of the muon collider to 10 TeV is able to probe heavy Higgs above the TeV scale. In panel (b) of Figure 7, we show the cross section of $H\to\mu^{\pm}\mu^{\pm}JJ$ as a function of the center-of-mass energy $\sqrt{s}$, which clearly depicts the logarithmic dependence of collision energy as in Equation 38.
@@
Here, the lower cut on $P_{T}(\mu)$ is selected heuristically. The best cut on the lower value of $P_{T}(\mu)$ indeed depends on $m_{N}$. We do not vary it in the analysis, because the difference of the significance is relatively small between the best and fixed cut on lower value of $P_{T}(\mu)$.
@@
Figure 9: The significance (upper panels) and precision (lower panels) of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature. In the left panels we have fixed the mass relation $m_{N}=m_{H}/3$. The blue, green, and red lines are the results for $\sin{\alpha}=0.15,0.1$, and $0.05$, respectively. In the right panels we have fixed $\sin{\alpha}=0.1$. The blue, green, and red lines are the results for $m_{H}=$300 GeV, 500 GeV, 700GeV. The dashed (solid) lines are for 3 (10) TeV muon collider with an integrated luminosity of $L=1(10)~\text{ab}^{-1}$.
@@
Based on the selection cuts in Table 2, we then explore the significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature at the muon collider. The results are shown in Figure 9. In the left-upper panel, we show the significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature as a function of $m_{N}$ for $\sin\alpha=0.15,0.10$, and 0.05 respectively, where we have fixed the mass relation $m_{N}=m_{H}/3$. The significance decreases as $m_{N}$ increases. On the 3TeV muon collider with $1~\text{ab}^{-1}$ integrated luminosity, we can reach 5 $\sigma$ discovery for $m_{N}\lesssim 330$ GeV and $\sin\alpha=0.15$. Lowering $\sin\alpha$ to 0.05, the significance still exceeds 5 $\sigma$ when $m_{N}\lesssim 200$ GeV. Increasing the collision energy to 10 TeV with 10 ab-1 data could extend the discovery limit up to $m_{N}\lesssim 1$ TeV with $\sin\alpha=0.1$. The precision at 3 TeV is about 10% $\sim$ 20% when $m_{N}\lesssim 300$ GeV, and it is less than 10% at 10 TeV when $m_{N}\lesssim 800$ GeV and $\sin\alpha\gtrsim 0.1$.
@@
In the right-upper panel of Figure 9, we show the significance of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature as a function of $m_{N}$ for $m_{H}=$300 GeV, 500 GeV and 700 GeV respectively, where we have fixed the Higgs mixing $\sin\alpha=0.1$. For $m_{H}=300$ GeV, the sensitivity reaches the maximum value of 23 $\sigma$ around $m_{N}=110$ GeV at the 3 TeV muon collider, and such result can reach to 130 $\sigma$ around $m_{N}=110$ GeV at the 10 TeV muon collider,. Meanwhile, for $m_{H}=500$ GeV and 700 GeV, we find that there exists a local minimum value of significance around $m_{N}\sim 135$ GeV. And the 5 $\sigma$ discovery could be realized when $m_{N}\lesssim 230$ GeV with $m_{H}=500$ GeV or $m_{N}\lesssim 320$ GeV with $m_{H}=700$ GeV at the 3 TeV muon collider, and $m_{N}\lesssim 240$ GeV with $m_{H}=500$ GeV or $m_{N}\lesssim 340$ GeV with $m_{H}=700$ GeV at the 10 TeV muon collider. Therefore, the 10 TeV muon collider is expected to probe most regions below the kinematic threshold $m_{N}\lesssim m_{H}/2$. We find a precision of $\mathcal{O}(10)\%$ at the 3 TeV muon collider and $\mathcal{O}(3)\%$ at the 10 TeV muon collider, respectively.
Figure 10: The discovery reach of the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature. In the left panel we have set $m_{N}=m_{H}/3$. The red region is excluded by direct searches of heavy Higgs Lane:2024vur . In the right panel, the red and green lines are the results for $\sin\alpha=0.1$ and $0.05$.
@@
The $5\sigma$ discovery region in the $m_{N}-\sin\alpha$ plane with $m_{N}=m_{H}/3$ is shown in the left panel of Figure 10. At the 3 TeV muon collider with $1~\text{ab}^{-1}$ data, the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature can be detected with the Higgs mixing parameter as small as $\sin\alpha\simeq 0.03$. According to the results in Figure 7, the cross section of $H\to\mu^{\pm}\mu^{\pm}JJ$ is heavily suppressed when $\sin\alpha\lesssim 0.01$, thus it becomes less promising. Due to the larger cross section and integrated luminosity, the 10 TeV muon collider could discover the Higgs mixing parameter as small as $\sin\alpha\simeq 0.005$ for heavy Higgs around the electroweak scale. With larger singlet scalar mass, the value of Higgs mixing parameter $\sin\alpha$ increases for 5 $\sigma$ discovery. Typically, for $\sin\alpha=0.15$, we might probe $m_{N}\lesssim 330$ GeV and $m_{N}\lesssim 950$ GeV for 3 TeV muon collider and 10 TeV muon collider separately.
@@
In the right panel of Figure 10, we show the sensitivity region in the $m_{H}-m_{N}$ plane with fixed $\sin\alpha$ of 0.1 and 0.05. At the 3 TeV muon collider, a large parameter space can be detected for $230~{\rm{GeV}}<m_{H}<880~\rm{GeV}$ and $m_{N}<340~\rm{GeV}$ with $\sin\alpha=0.1$. Of course, the sensitive region heavily depends on the Higgs mixing parameter. For $\sin\alpha>0.1$, although certain region with $m_{H}\lesssim 600$ GeV is already excluded by current direct searches of heavy Higgs, $m_{H}$ around TeV scale is also promising with such large $\sin\alpha$. When $\sin\alpha=0.05$, the promising region becomes $230~{\rm{GeV}}<m_{H}<640~\rm{GeV}$ and $m_{N}<220~\rm{GeV}$, which is clearly smaller than the scenario with $\sin\alpha=0.1$. Increasing the collision energy to 10 TeV will greatly enlarge the discovery region. For instance, the parameter space can be detected for $220~{\rm{GeV}}<m_{H}<2800~\rm{GeV}$ and $m_{N}<1100~\rm{GeV}$ with $\sin\alpha=0.1$, as well as $220~{\rm{GeV}}<m_{H}<2000~\rm{GeV}$ and $m_{N}<800~\rm{GeV}$ for $\sin\alpha=0.05$.